25 June 2000  Page 1 of 1  J3/00-243

Subject: Issues 277-279 From: Van Snyder

1 Introduction

Issues 277-279 are concerned with problems in the discussion of main program arguments.

2 Edits

Edits refer to 00-007r2. Page and line numbers are displayed in the margin. Absent other instructions, a page and line number or line number range implies all of the indicated text is to be replaced by immediately following text, while a page and line number followed by + indicates that immediately following text is to be inserted after the indicated line. Remarks for the editor are noted in the margin, or appear between [ and ] in the text.

[Editor: Delete issue 277.] 236:19-25

If a main program argument is a default character scalar it is the command program argument. 236:28-29 It shall have assumed length.

If a main program argument is a default character array of rank one, it is the argument text 236:33-34 program argument. It shall have assumed length, assumed shape, and an explicit lower bound of zero. [Editor: Delete issue 278.] 236:35-42

[Editor: Replace “The processor ... program argument” by “The processor shall assign to the 237:1-2 argument text program argument a representation of the command name and the command arguments”.]

[Editor: Delete issue 279.] 237:4-10

If a main program argument is a default integer array of rank one it is the argument length 237:14-15 program argument. It shall have assumed shape, and an explicit lower bound of zero.